iol_equipment_downtime_maintenance_log
 Share
The version of the browser you are using is no longer supported. Please upgrade to a supported browser.Dismiss

 
$
%
123
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
 
|
 
Still loading...
ABCDEF
1
PHOIOL: 5th Floor Integrated Optics Lab Equipment Maintenance Downtime Log
2
3
Color Code DefinitionsDOWN - Equipment not usable
4
UP - Equipment is operational
5
Partially Operational - Equipment may be usable contact Lab Manager
6
7
DateEquipmentStatusFollow on NotesSystem Update
8
9
10
12/09/10Bruker FTIRUPReplaced the HeNe and set up the new computer.Service Complete
11
12/03/10Bruker FTIRDOWNHeNe needs replacement. Service has been called. Service Call Completed
12
07/27/10Bruker FTIRUPRealigned system motors and optics. System is working great!Service Complete
13
07/15/10Varian Cary 5000UPReadjusted tracking of beam. System is working great!Service Complete
14
07/12/10Bruker FTIRDOWNAlignment is needed. Service has been called. Service Call Completed
15
07/12/10Varian Cary 5000DOWNEquipment is down. Service will be contacted.Service Call Completed
16
01/22/10Bruker FTIRUPService was in and alignments were done. System is scanning.Service Complete
17
01/20/10FC150DOWNStage and thermode(thermocouple) are fixed. Temp controller has been sent out for repair. Will be back first week in February.Service in process
18
01/13/10Bruker FTIRDOWNFTIR is not scanning. It is DOWN. Service will be contacted.Service in process
19
01/12/10FC150DOWNThermode (thermocouple) and stage has been sent for replacement. Service will reinstall it on Tuesday and hopefully this will fix the issue with the temp controller not working correctly.Service in process
20
12/08/09FC150DOWNTemp controller is defaulting parameters. Looking to replace them with new type.Service in process
21
09/09/09Varian Cary 5000UPChopper was replaced. It is back up and running.Service Complete
22
08/28/09Varian Cary 5000DOWNChopper needs to be replaced. Parts on order. Back up in 2 weeks.Parts on Order
23
02/04/09FC150UPUsing borrowed temp controller till we receive our back. Equipment is up and running.Service Complete
24
01/26/09Bruker FTIRUPThe alignment has been done. FTIR is back up and running.Service Complete
25
01/17/09Bruker FTIRDOWNService has been called and will arrive on Wednesday, 1-22-09Service in process
26
11/19/08FC150UPTemp controller has been fixed and put back into the system. Will get PM in a few weeks.Service Complete
27
11/04/08FC150DOWNWaiting for temp controller to be fixed and finish PM. Equipment will be down for up to 2 weeks.Part Being Fixed
28
10/31/08FC150DOWNService will be arriving Tuesday, 11-4-08. Should be up and running soon after.Service in process
29
10/27/08FC150DOWNGetting multiple errors.Service Call Completed
30
09/20/08plasma cleanerUPFixed hose connection.Service Complete
31
09/17/08plasma cleanerDOWNDoes not seem to be working.Service in process
32
03/14/08FC150UPCoupling part replacement complete.Service Complete
33
03/12/08FC150DOWNWe had the wrong part needed. The correct part is on order. Not sure on the lead time yet.Parts on Order
34
03/11/08FC150DOWNCoupler will be replaced on Wednesday 3/12/08.Service in process
35
03/05/08FC150DOWNCoupler needs to be replaced on motor.Service Call Completed
36
12/20/07FC150UPService CompleteService Complete
37
12/17/07FC150DOWNThermal controllers are blinking, recalibration procedure did not work - service contacted and coming in on Thursday December 20thService in process
38
10/15/07FC150UPPM CompleteMaintenance Complete
39
10/11/07FC150DOWNTechnician from Suss Micro-tech will be performing standard PM on the system Thursday and FridayStandard Maintenance
40
09/28/07House VacuumUPPM CompleteMaintenance Complete
41
09/27/07House VacuumDOWNOil Filter Change on Vacuum PumpStandard Maintenance
Loading...
 
 
 
Equipment Maintenance Log
Listings
Sheet3